The second angle of a triangle is 3times the first and third is the 12 less than twice the first . Find the measure if the angles .
Answer by aswathytony(47)

sum of the angles of a triangle is 180deg.
let the angle measure of 1st, 2nd & 3rd angles are x, y & z respectively.
given second angle is three times first i.e y = 3 * x = 3x ......(1)
third angle is 12 less than twice the first
i.e z = 2* x - 12 = 2x - 12 ............... (2)
also x + y + z = 180 .........(3)
substituting (1) & (2) in (3)
x + 3x + 2x - 12 = 180
6x - 12 = 180
6x = 180 + 12 = 192
x = 192 / 6 = 32deg

y = 3x = 3 * 32 = 96 deg
z = 2x - 12 = 2 * 32 - 12 = 64 -12 = 52 deg